OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (2, 16) A. y = x B. y = x + 6 C. y = x – 6 D. y = x + 6

OpenStudy (jdoe0001):

what's the slope of " y = x + 8 "?

OpenStudy (anonymous):

parallel lines share the same slope perpendicular lines' slopes are negative reciprocals of each other memorize that

OpenStudy (texaschic101):

y = x + 8 slope is 1 y - y1 = m(x - x1) using slope 1 and points (2,16) y - 16 = 1(x - 2) y - 16 = x - 2 y = x - 2 + 16 y = x + 14 Did you write the problem correctly ?

OpenStudy (anonymous):

i guess not still having problems understanding the problem :(